∠A and ∠B are supplementary angles. m ∠A is 9° less than twice m ∠B. Find m ∠A and m ∠B.
Answer by TimothyLamb(4379)

x = measure of angle A (degrees)
y = measure of angle B (degrees)
---
x + y = 180
x = 2y - 9
---
put the system of linear equations into standard form
---
x + y = 180
x - 2y = -9
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = 117
y = 63
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---